OpenStudy (haleyelizabeth2017):

If sin theta=3/5 and theta has its terminal side in Quadrant II, find the exact value of tan 2 theta.

OpenStudy (haleyelizabeth2017):

So would this just be tan(2 theta)=2*sin(theta)*tan(theta)?

OpenStudy (haleyelizabeth2017):

Or am I going about this wrong?

jimthompson5910 (jim_thompson5910):

if it were sin(2theta), then you'd be correct

jimthompson5910 (jim_thompson5910):

look at this pdf http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

jimthompson5910 (jim_thompson5910):

look at the `Double Angle Formulas` on page 2

OpenStudy (haleyelizabeth2017):

Oh LOL

OpenStudy (haleyelizabeth2017):

I know that tan would be 3/4...So, I just plug that in for theta?

jimthompson5910 (jim_thompson5910):

you replace every `tan(theta)` with 3/4

OpenStudy (haleyelizabeth2017):

I'm confused...so the numerator would just be 2(3/4)?

OpenStudy (haleyelizabeth2017):

or 2tan(3/4)?

jimthompson5910 (jim_thompson5910):

\[\Large \tan(2\theta) = \frac{2\tan(\theta)}{1-\tan^2(\theta)}\] \[\Large \tan(2\theta) = \frac{2\tan(\theta)}{1-(\tan(\theta))^2}\] \[\Large \tan(2\theta) = \frac{2*({\color{red}{\tan(\theta)}})}{1-({\color{red}{\tan(\theta)}})^2}\] \[\Large \tan(2\theta) = \frac{2*({\color{red}{3/4}})}{1-({\color{red}{3/4}})^2}\] \[\Large \tan(2\theta) = ???\]

OpenStudy (haleyelizabeth2017):

Ohhh

OpenStudy (haleyelizabeth2017):

(3/2)/(1-(9/16)) right?

jimthompson5910 (jim_thompson5910):

yes, now simplify as much as possible

OpenStudy (haleyelizabeth2017):

Okay

OpenStudy (haleyelizabeth2017):

24/7?

jimthompson5910 (jim_thompson5910):

I'm getting the same

jimthompson5910 (jim_thompson5910):

actually hold on

jimthompson5910 (jim_thompson5910):

I'm just realizing that theta is in Q2

jimthompson5910 (jim_thompson5910):

tangent is negative in Q2

OpenStudy (haleyelizabeth2017):

so -24/7? LOL

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (haleyelizabeth2017):

Awesome, again, thank you for the help!

jimthompson5910 (jim_thompson5910):

glad to be of help

OpenStudy (haleyelizabeth2017):

:)